OpenStudy (helpblahblahblah):

Medal plz help! :) Which calculation can be used to find the value of q in the equation q3 = 64? A. q equals square root of 64 over 3 B. q equals square root of 64 C. q equals cube root of 64 D. q equals 3 multiplied by square root of 64

hartnn (hartnn):

\(q^3 = 64\) take cube root on both sides

hartnn (hartnn):

\(\sqrt[3] {q^3} = q\)

hartnn (hartnn):

so left side, you have 'q' right side you have cube root of 64

OpenStudy (helpblahblahblah):

so C

hartnn (hartnn):

yes :)

OpenStudy (helpblahblahblah):

thxx! :D